Grade 12th passMechanics

If two different projectiles have same range and time periodsT1 AndT2 then T1 T2 is proportional to

akhil
10 Years agoGrade 12th pass

1 Answer

Vikas TU
10 Years ago
same Range 
: – >
.i.e. u^2sin2@/g = v^2sin2#/g
or
u^2sin2@ = v^2sin2# …...........(1)
 
T1 = 2usin@/g  and   T2 =2vsin#/g
T1*T2 = 4uvsin@sin#/g^2